# On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective

### v1

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have recently emerged as a novel approach to function approximation, demonstrating remarkable potential in various domains. Despite their theoretical promise, the robustness of KANs under adversarial conditions has yet to be thoroughly examined. In this paper, we explore the adversarial robustness of KANs, with a particular focus on image classification tasks. We assess the performance of KANs against standard white-box adversarial attacks, comparing their resilience to that of established neural network architectures. Further, we investigate the transferability of adversarial examples between KANs and Multilayer Perceptron (MLPs), deriving critical insights into the unique vulnerabilities of KANs. Our experiments use the MNIST, FashionMNIST, and KMNIST datasets, providing a comprehensive evaluation of KANs in adversarial scenarios. This work offers the first in-depth analysis of security in KANs, laying the groundwork for future research in this emerging field.
@@ 1 Introduction
In the rapidly evolving field of deep learning, the robustness of neural networks against adversarial attacks has emerged as a cornerstone of research, driven by the ubiquity of their application, often in sensitive areas. As a result, a multitude of methods have been developed to detect and mitigate adversarial attacks, underscoring the critical need for robust defenses in real-world scenarios. These methods range from adversarial training [31, 4, 39, 2] and defensive distillation [36, 35, 6] to approaches like feature squeezing [52], input transformations [17, 18], and the use of auxiliary models for detection purposes [55, 37, 27]. The proliferation of these techniques reflects the ongoing efforts within the research community to fortify neural networks against increasingly sophisticated attacks.
@@
Multilayer Perceptrons (MLPs)—a foundational model in deep learning—have been extensively studied for their vulnerability and defense mechanisms against adversarial attacks [12, 45, 33, 21, 20]. Recently, the introduction of Kolmogorov-Arnold Networks (KANs) [29, 5, 15, 43, 22] has opened new avenues in function approximation, with theoretical promises of enhanced performance and efficiency.
@@
Many studies have compared the performance of MLPs and KANs across a range of tasks. For example, Yu et al. [53] found that MLPs generally perform better in areas such as machine learning and computer vision. Conversely, KANs excel in representing symbolic formulas.
@@
Another interesting comparison was done by Zeng et al. [54], focusing on the performance of KANs and MLPs in handling irregular or noisy functions. They found that KANs exhibited superior performance both for regular and irregular functions. However, in cases involving functions with jump discontinuities or singularities, MLPs demonstrated greater efficacy.
@@
Recently, Dong et al. [13] studied KANs for time-series classification. They concluded that KANs exhibit significant robustness advantages, attributed to their lower Lipschitz constants. Note that they focused on time-series problems, while we focus herein on image tasks; further, they deployed a single attack algorithm (Projected Gradient Descent, PGD [31]), whereas we will deploy four attack algorithms (including PGD).
@@
This paper delves into the robustness or KANs and MLPs when faced with adversarial attacks under white-box conditions, and investigates the transferability of adversarial examples between these two architectures. By examining how adversarial examples crafted for one model type might affect another, we aim to uncover unique susceptibilities associated with KANs. Using the MNIST, KMNIST, and FashionMNIST datasets, our analysis not only sheds light on the robustness of KANs but also contributes to the broader understanding of neural network security in adversarial settings.
In the next section we provide some background on MLPs, KANs, and robustness. Section 3 studies the robustness of KANs against white-box attacks. Section 4 peruses the question: How transferable are adversarial examples created by KANs? Finally, we end with concluding remarks in Section 5.
@@ 2 Background
2.1 Multilayer Perceptrons
Multilayer Perceptrons (MLPs) are a class of feedforward artificial neural networks (ANNs) that consist of multiple layers of nodes, typically organized into an input layer, one or more hidden layers, and an output layer. Each node in one layer is connected to every node in the subsequent layer, with weights assigned with these connections, which are adjusted during the training process. MLPs are capable of approximating complex functions and have been widely used in various fields, such as image recognition, speech analysis, and natural language processing.
Mathematically, let $\mathbf{x}\in\mathbb{R}^{d}$ represent the input vector. Define the output of the $i$-th layer as $\mathbf{h}_{i}$, with $\mathbf{h}_{0}=\mathbf{x}$ for the input layer. Let $\mathbf{W}_{i}\in\mathbb{R}^{n_{i}\times n_{i-1}}$ and $\mathbf{b}_{i}\in\mathbb{R}^{n_{i}}$ be the weight matrix and bias vector for the $i$-th layer, respectively, where $n_{i}$ denotes the number of nodes in layer $i$. The output of the $i$-th layer, $\mathbf{h}_{i}$, can be expressed as:
$\mathbf{h}_{i}=\sigma(\mathbf{W}_{i}\mathbf{h}_{i-1}+\mathbf{b}_{i}),$
where $\sigma(\cdot)$ denotes the activation function, which is often ReLU [34] (Rectified Linear Unit) or sigmoid. The final output of the MLP, denoted as $\mathbf{y}$, is given by the output of the last layer $L$:
$\mathbf{y}=\mathbf{h}_{L}=\sigma(\mathbf{W}_{L}\mathbf{h}_{L-1}+\mathbf{b}_{L}).$
Training an MLP involves optimizing the weights $\{\mathbf{W}_{i}\}$ and biases $\{\mathbf{b}_{i}\}$ to minimize a loss function, typically using gradient-based methods, such as backpropagation.
The universal approximation theorem, established by Cybenko [9] and Hornik et al. [19], states that an MLP with a single hidden layer and a finite number of neurons can approximate any continuous function on a compact subset of $\mathbb{R}^{n}$ with suitable weights and activation functions.
2.2 Kolmogorov-Arnold Networks
Kolmogorov Arnold Networks (KANs) are based on the Kolmogorov-Arnold representation theorem [23], which asserts that any multivariate continuous function can be represented as a finite superposition of univariate functions and additions. KANs leverage this theorem to construct neural networks with a specific architecture that is different from standard MLPs.
Mathematically, given a continuous function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the Kolmogorov-Arnold theorem states that there exist continuous functions $\{\varphi_{i}\}$ and $\{\psi_{ij}\}$ such that:
$f(\mathbf{x})=\sum_{i=1}^{2d+1}\varphi_{i}\left(\sum_{j=1}^{d}\psi_{ij}(x_{j})\right).$
In the context of KANs, the $\varphi_{i}$ functions correspond to the activation functions applied at the hidden nodes, and the $\psi_{ij}$ functions represent the preprocessing applied to the inputs. This leads to a network architecture where the input is first transformed by a set of univariate functions, followed by linear combinations and further nonlinear transformations.
KANs offer a different approach to function approximation compared to MLPs, potentially providing advantages in terms of expressiveness and robustness under certain conditions.
KANs leverage the aforementioned theorem to construct neural-network architectures that decompose complex functions into simpler, univariate components. This approach contrasts with the layer-by-layer processing in MLPs, offering potential advantages in terms of interpretability and parameter efficiency.
The structure of KANs suggests that they might be inherently more robust to certain types of perturbations compared to traditional MLPs, as the decomposition into univariate functions can isolate and mitigate the impact of input variations.
2.3 Robustness in Artificial Neural Networks
@@
Robustness in artificial neural networks (ANNs) refers to their ability to maintain performance when subjected to various forms of perturbations, such as noise, adversarial attacks, or distributional shifts. A robust ANN should not only perform well on the training and validation datasets but also generalize effectively to unseen data, including those with slight modifications or corruptions.
@@
Mathematically, robustness can be defined as follows: Consider an ANN $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, parameterized by $\theta$, and let $\mathbf{x}\in\mathbb{R}^{n}$ be an input to the network. The network is said to be robust to perturbations if, for a small perturbation $\mathbf{\delta}\in\mathbb{R}^{n}$, the output of the network remains unchanged, that is:
@@
Adversarial robustness in particular has gained significant attention in recent years. Adversarial attacks involve making small—often imperceptible—perturbations to input data that can cause a neural network to produce incorrect outputs with high confidence [16]. These attacks exploit the vulnerability of neural networks to specific input manipulations, raising concerns about their reliability in real-world applications [41].
@@
For example, the Fast Gradient Sign Method (FGSM) demonstrates how slight changes to image pixels can deceive a classifier into misclassification [16]. Additionally, the Carlini-Wagner (C&W) attack shows that even robust defenses can be circumvented through optimized perturbations [7].
@@
Such examples underscore the critical need for developing networks with enhanced robustness against adversarial attacks. Numerous other attack methodologies further highlight this necessity, revealing the extensive landscape of adversarial vulnerability in neural networks [16, 28, 41, 42, 48, 26, 8, 7, 25, 3, 24].
@@
Several techniques have been proposed to enhance the robustness of neural networks, including adversarial training, regularization methods, and architectural modifications. Adversarial training involves training the network on adversarially perturbed examples, thereby improving its ability to withstand attacks [31]. Regularization methods—such as dropout [40] and weight decay—help prevent overfitting and improve generalization. Architectural modifications—such as incorporating attention mechanisms [44] or leveraging alternative network structures—also contribute to robustness.
@@
The upshot is that robustness is a critical aspect of ANN performance, particularly in the context of adversarial attacks and other perturbations. By understanding and improving the robustness of ANNs, we can develop more reliable and secure models for various applications.
3 White-Box Attacks on KANs
@@
Do KANs offer enhanced security compared to MLPs when confronted with a white-box adversary?
@@
To address this question we conduct an experimental evaluation of KANs. While it may appear unconventional to begin with experiments, direct assessment of the security of KANs is most effectively achieved through targeted attacks and subsequent analyses. We begin our study of KAN robustness with a white-box adversary because it represents the most formidable threat model.
3.1 Threat Model
We consider an adversary with full knowledge of the model’s architecture and trained parameters. We assume that the adversary can perturb the original input $\mathbf{x}$ to generate an adversarial example $\mathbf{x_{\text{adv}}}$ within a specified perturbation bound, such that $\|\mathbf{x}-\mathbf{x_{\text{adv}}}\|_{\infty}\leq\epsilon$. For all the datasets, we set $\epsilon=32/255$.
The adversary is considered successful if it can generate an adversarial input $\mathbf{x_{\text{adv}}}$ within this bound that the classifier misclassifies. This scenario is known as an untargeted attack.
Formally, given a classifier $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ and an original input $\mathbf{x}$ with ground truth label $y$, the goal of the untargeted adversary is to find an adversarial example $\mathbf{x_{\text{adv}}}$ such that:
maximize $\mathcal{L}(f(\mathbf{x_{\text{adv}}}),y)$
subject to $\|\mathbf{x_{\text{adv}}}-\mathbf{x}\|_{\infty}\leq\epsilon,$
where $\mathcal{L}(\cdot)$ is the loss function used to train the classifier. The adversarial example $\mathbf{x_{\text{adv}}}$ is successful if the classifier’s prediction on $\mathbf{x_{\text{adv}}}$ differs from the true label $y$, i.e., $f(\mathbf{x_{\text{adv}}})\neq y$. This formulation encapsulates the adversary’s objective of inducing a misclassification while adhering to the perturbation constraints.
To evaluate a model’s robustness, we first select a set of clean test examples that are correctly classified by the model. We then generate adversarial examples from this set using one of four different attacks. The security of the model is quantified by the percentage of examples that remain correctly classified under these adversarial conditions. Given the relatively recent introduction of KANs, we explore a broad range of attacks and models in our experiments. Below, we enumerate the attacks and models utilized in this study, along with the rationale for their inclusion.
3.2 White-Box Attacks
We evaluate the robustness of our models against four distinct white-box attacks. We begin with the Fast Gradient Sign Method (FGSM) [16], a foundational attack that serves as an initial test of model resilience. Subsequently, we assess the models using more potent iterative attacks, including the Projected Gradient Descent (PGD) [31], Carlini & Wagner (C&W) [7], and Momentum Iterative Method (MIM) [14]. The hyperparameters used for these attacks are outlined in Table 1. A detailed description of each attack is provided in Appendix B.
Table 1: Hyperparameters employed in experiments with white-box attacks.
Attack Parameters
FGSM $\epsilon=32/255$
PGD $\epsilon=32/255$, steps = 30, $\alpha=0.01$
C&W $\epsilon=32/255$, steps = 30, $\alpha=0.01$, $c=10$
MIM $\epsilon=32/255$, steps = 30, $\alpha=0.01$
@@ 3.3 Classifier Models
We used six different neural network models, comprising three KAN architectures and three MLP architectures. These models were categorized into small, medium, and large configurations, as specified in Table 2. The KAN models were configured with a uniform setting of num knots = 5 and spline order = 3 across all sizes. Each model was trained on the MNIST [11], FashionMNIST [50], and KMNIST [38] datasets for 20 epochs. Training was performed using the AdamW [30] optimizer, with learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-4}$, and batch size 64.
Table 2: Model configurations.
@@
3.4 Analysis of White-Box Attacks
Table 3 shows the results of our experiments of white-box attacks on KANs and MLPs.

### v2

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have recently emerged as a novel approach to function approximation, demonstrating remarkable potential in various domains. Despite their theoretical promise, the robustness of KANs under adversarial conditions has yet to be thoroughly examined. In this paper we explore the adversarial robustness of KANs, with a particular focus on image classification tasks. We assess the performance of KANs against standard white-box and black-box adversarial attacks, comparing their resilience to that of established neural network architectures. Our experimental evaluation encompasses a variety of standard image classification benchmark datasets and investigates both fully connected and convolutional neural network architectures, of three sizes: small, medium, and large. We conclude that small- and medium-sized KANs (either fully connected or convolutional) are not consistently more robust than their standard counterparts, but that large-sized KANs are, by and large, more robust. This comprehensive evaluation of KANs in adversarial scenarios offers the first in-depth analysis of KAN security, laying the groundwork for future research in this emerging field.
@@ 1 Introduction
In the rapidly evolving field of deep learning, the robustness of neural networks against adversarial attacks has emerged as a cornerstone of research, driven by the ubiquity of their application, often in sensitive areas. As a result, a multitude of methods have been developed to detect and mitigate adversarial attacks, underscoring the critical need for robust defenses in real-world scenarios. These methods span several areas, including adversarial training (Madry et al., 2017; Bai et al., 2021; Shafahi et al., 2019; Andriushchenko & Flammarion, 2020), defensive distillation (Papernot et al., 2016; Papernot & McDaniel, 2016; Carlini & Wagner, 2016), feature squeezing (Xu et al., 2017), input transformations (Guo et al., 2017; Harder et al., 2021), and using auxiliary models for detection purposes (Zheng & Hong, 2018; Pinhasov et al., 2024; Lapid et al., 2024a). The proliferation of these techniques reflects the ongoing efforts within the research community to fortify neural networks against increasingly sophisticated attacks.
@@
Fully connected neural networks (FCNNs) and convolutional neural networks (CNNs)—foundational models in deep learning—have been extensively studied for their vulnerability and defense mechanisms against adversarial attacks (Dey et al., 2017; Wang et al., 2020; Mohammed et al., 2020; Kerlirzin & Vallet, 1993; Kalina et al., 2022). Recently, the introduction of Kolmogorov-Arnold Networks (KANs) has opened new avenues in function approximation, with theoretical promises of enhanced performance and efficiency (Liu et al., 2024; Bozorgasl & Chen, 2024; Genet & Inzirillo, 2024; Vaca-Rubio et al., 2024; Kiamari et al., 2024).
@@
Many studies have compared the performance of KANs across a range of tasks. For example, Yu et al. (2024) found that FCNNs generally perform better in areas such as machine learning and computer vision. Conversely, KANs excel in representing symbolic formulas.
@@
Another interesting comparison was done by Zeng et al. (2024), focusing on the performance of KANs and FCNNs in handling irregular or noisy functions. They found that KANs exhibited superior performance both for regular and irregular functions. However, in cases involving functions with jump discontinuities or singularities, FCNNs demonstrated greater efficacy.
@@
Dong et al. (2024) studied KANs for time-series classification. They concluded that KANs exhibit significant robustness advantages, attributed to their lower Lipschitz constants. Note that they focused on time-series problems, while we focus herein on image-related tasks; further, they deployed a single attack algorithm (Projected Gradient Descent, PGD; Madry et al. (2017)), whereas we will deploy seven attack algorithms.
@@
This paper delves into the robustness of KANs when faced with adversarial attacks under white-box and black-box conditions. Using the MNIST, KMNIST, FashionMNIST, CIFAR-10, and SVHN datasets, our analysis not only sheds light on the robustness of KANs but also contributes to the broader understanding of neural network security in adversarial settings.
The next section provides a comprehensive overview of the architectures employed in this study, accompanied by a detailed discussion of their robustness. Section 3 outlines the experimental setup used to evaluate our proposed approach. Section 4 examines the robustness of the different architectures against adversarial attacks. Finally, we present concluding remarks in Section 5.
@@ 2 Background
2.1 Model Architectures
Our study evaluates two primary types of layers used in artificial neural networks: fully connected layers and convolutional layers. Fully connected layers are the foundational building blocks of feedforward neural networks, where each neuron is connected to every neuron in the next layer. FCNNs are classical feedforward networks that learn floating-point weights for these connections (Cybenko, 1989). In contrast, FCKANs, inspired by the Kolmogorov-Arnold theorem (Kolmogorov, 1961), decompose functions into simpler univariate components, learning these components instead of conventional weights, as detailed by Liu et al. (2024).
Convolutional layers extend the capabilities of fully connected layers by leveraging spatial hierarchies in data such as images. These layers are the first and main layers introduced in CNNs (LeCun et al., 1998), where they typically learn floating-point weights for their convolutional filters. CKANs (Drokin, 2024) apply their unique decomposition principle to learn univariate functions for the filters rather than conventional weights. For more detailed descriptions, please refer to the cited literature.
2.2 Robustness in Neural Networks
@@
Robustness in neural networks refers to their ability to maintain performance when subjected to various forms of perturbations, such as noise, adversarial attacks, or distributional shifts. A robust neural network should not only perform well on the training and validation datasets but also generalize effectively to unseen data, including those with slight modifications or corruptions.
@@
Mathematically, robustness can be defined as follows: Consider a neural network $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, parameterized by $\theta$, and let $\mathbf{x}\in\mathbb{R}^{n}$ be an input to the network. The network is said to be robust to perturbations if, for a small perturbation, $\mathbf{\delta}\in\mathbb{R}^{n}$, the output of the network remains unchanged, that is:
@@
Adversarial robustness in particular has gained significant attention in recent years. Adversarial attacks involve making small—often imperceptible—perturbations to input data that can cause a neural network to produce incorrect outputs with high confidence (Goodfellow et al., 2014). These attacks exploit the vulnerability of neural networks to specific input manipulations, raising concerns about their reliability in real-world applications (Szegedy et al., 2013).
@@
For example, the Fast Gradient Sign Method (FGSM) demonstrates how slight changes to image pixels can deceive a classifier into misclassification (Goodfellow et al., 2014). Additionally, the Carlini-Wagner (C&W) attack shows that even robust defenses can be circumvented through optimized perturbations (Carlini & Wagner, 2017).
@@
Such examples underscore the critical need for developing networks with enhanced robustness against adversarial attacks. Numerous other attack methodologies further highlight this necessity, revealing the extensive landscape of adversarial vulnerability in neural networks (Goodfellow et al., 2014; Lapid et al., 2024b; Szegedy et al., 2013; Tamam et al., 2023; Wei et al., 2022; Lapid et al., 2022; Chen et al., 2024; Carlini & Wagner, 2017; Lapid & Sipper, 2023b; Andriushchenko et al., 2020; Lapid & Sipper, 2023a).
@@
Several techniques have been proposed to enhance the robustness of neural networks, including adversarial training, regularization methods, and architectural modifications. Adversarial training involves training the network on adversarially perturbed examples, thereby improving its ability to withstand attacks (Madry et al., 2017). Regularization methods—such as dropout (Srivastava et al., 2014) and weight decay—help prevent overfitting and improve generalization. Architectural modifications—such as incorporating attention mechanisms (Vaswani et al., 2017) and leveraging alternative network structures—also contribute to robustness.
@@
The upshot is that robustness is a critical aspect of neural network performance, particularly in the context of adversarial attacks and other perturbations. By understanding and improving the robustness of neural networks, we can develop more reliable and secure models for various applications.
2.3 Adversarial Attack Types
Adversarial attacks are techniques used to evaluate the robustness of neural networks by introducing small perturbations to input data, causing the model to misclassify the input. These attacks are broadly categorized into white-box attacks and black-box attacks, based on the adversary’s level of access to the model.
White-Box Attacks. In white-box attacks, the adversary has full knowledge of the target model, including its architecture, weights, and gradients. This access allows the adversary to compute perturbations directly by exploiting the model’s vulnerabilities. White-box attacks represent the strongest adversarial threat model, as they assume complete transparency of the system.
Black-Box Attacks. In black-box attacks, the adversary has no direct access to the model’s internal parameters or gradients. Instead, the adversary generates adversarial examples by querying the model or leveraging transferability across architectures. Black-box attacks are more practical in real-world scenarios, as they mimic situations where the adversary does not have access to proprietary models or data.
3 Experimental Setup
@@
Do KANs offer enhanced security compared to classical neural network architectures, when confronted with an adversary?
@@
To address this question we conduct an experimental evaluation of KANs. While it may appear unconventional to begin with experiments, direct assessment of the security of KANs is most effectively achieved through targeted attacks and subsequent analyses. We evaluate the robustness of FCKANs, CKANs, FCNNs and CNNs under white-box and black-box adversarial attacks. Both fully connected and convolutional architectures are tested across multiple datasets, and the results are analyzed to compare the robustness of the two architectures.
@@ 3.1 Classifier Models
We utilized six models for each architecture (fully connected and convolutional), categorized into small, medium, and large configurations. The configurations for fully connected models are detailed in Table 1, while those for convolutional models are provided in Table 2. All the KAN models were configured with a uniform setting of num knots = 5 and spline order = 3.
Table 1: Model configurations for fully connected architectures.
@@
Table 2: Model configurations for convolutional architectures. All convolutional layers use a kernel size of $3\times 3$, stride of 1, and padding of 1. Max pooling is applied after the last convolutional layer, followed by two fully connected layers with hidden dimensions [1024, 256].
(a) Grayscale Datasets
Model Output Channels #Params
[…]
(b) Multi-Channel Datasets
Model Output Channels #Params
[…]
All the models were trained using the AdamW (Loshchilov & Hutter, 2017) optimizer for 20 epochs. The fully connected models were trained on the MNIST (Deng, 2012), FashionMNIST (Xiao et al., 2017), and KMNIST (Prabhu, 2019) datasets with a learning rate of $1\times 10^{-4}$, weight decay of $5\times 10^{-4}$, and a batch size of 64. The convolutional models were trained on the MNIST, SVHN (Netzer et al., 2011), and CIFAR-10 (Krizhevsky, 2009) datasets using a learning rate of $1\times 10^{-4}$, weight decay of $1\times 10^{-4}$, and a batch size of 32.
3.2 Adversarial Attacks
We evaluated the robustness of our models against four distinct white-box attacks. The evaluation began with FGSM (Goodfellow et al., 2014), a foundational attack used as an initial test of model resilience. Subsequently, we assessed the models using more advanced iterative attacks, including PGD (Madry et al., 2017), C&W (Carlini & Wagner, 2017), and the Momentum Iterative Method (MIM) (Dong et al., 2018).
Furthermore, we evaluated the robustness of all the models against three different black-box attack methodologies: transfer-based attacks (Section B.2.1), simple black-box attack (SimBA, (Tu et al., 2019)), square attack (Andriushchenko et al., 2020), and natural evolutionary strategies attack (NES, (Ilyas et al., 2018)). White-box attacks were evaluated on all test sets, while black-box attacks were evaluated on a random sample of 1,000 images from the corresponding test set. Transferability metrics were calculated using 5,000 images from the test sets.
The hyperparameter configurations for these experiments are provided in Table 3. For the grayscale datasets (MNIST, FashionMNIST, and KMNIST), an $\epsilon$ value of $32/255$ was utilized, whereas for the multi-channel; datasets (CIFAR-10 and SVHN), an $\epsilon$ value of $8/255$ was employed. All experiments were conducted with a batch size of 16.
Table 3: Hyperparameters used in experiments with white-box and black-box attacks.
Attack Type Attack Parameters (Grayscale Datasets) Parameters (Multi-channel Datasets)
White-box PGD $k=40$, $\alpha=0.01$ $k=30$, $\alpha=0.01$
C&W $k=40$, $\alpha=0.01$, $c=1$ $k=30$, $\alpha=0.01$, $c=1$
MIM $k=40$, $\alpha=0.01$, $\mu=1$ $k=30$, $\alpha=0.01$, $\mu=1$
Black-box Square $k=3000$ $k=2000$
SimBA $k=5000$ $k=3000$
NES $k=300$, $\alpha=0.025$, $\sigma=0.0625$, $n=40$ $k=200$, $\alpha=0.01$, $\sigma=0.01$, $n=30$
A comprehensive description of each attack can be found in Appendix B.
4 Results
4.1 Fully Connected Models
In the context of iterative white-box adversarial attack methods (PGD, C&W, MIM), small and medium FCKANs demonstrated lower robustness compared to small and medium FCNNs, as shown in Table 4. Across all examined datasets, small and medium FCKANs were notably vulnerable to these attacks. In contrast, small and medium FCNNs demonstrated a measurable—albeit limited—degree of robustness. Nonetheless, even these models remain highly susceptible to adversarial exploitation, rendering them far from secure in practice.